ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 130, 2002

Blasi, F. S. de and N. V. Zhivkov The number of nearest and farthest points to a compactum in Euclidean space

347

Caddeo, R., S. Montaldo and C. Oniciuc Biharmonic submanifolds in spheres

109

Castro, A. A. Backward inducing and exponential decay of correlations for partially hyperbolic attractors

29

Coven, E. M., W. Geller, S. Silberger and W. Thurston The symbolic dynamics of tiling the integers

21

Dolgopyat, D. On mixing properties of compact group extensions of hyperbolic systems

157

Edigarian, A. Pluripolar hulls and holomorphic coverings  77

Freiman, G. A. and B. L. Granovsky Asymptotic formula for a partition function of reversible coagulation-fragmentation processes

259

Geller, W. See E. M. Coven

Gordeev, N. and J. Saxl Products of conjugacy classes in Chevalley groups I. Extended covering numbers

207

Gordeev, N. and J. Saxl Products of conjugacy classes in Chevalley groups II. Covering and generation

249

Granovsky, B. L. See G. A. Freiman

Grundman, H. G., D. B. Leep and T. L. Smith ℚ-Adequacy of Galois 2-extensions

11

Gupta, S. K. and K. E. Hare Singularity of orbital measures in

SU(n) 93

Hare, K. E. See S. K. Gupta

Hoffman, C. and D. Rudolph A dyadic endomorphism which is Bernoulli but not standard

365

Josefson, B. Subspaces of l^{∞}(Γ) without quasicomplements  281

Leep, D. B., T. L. Smith and R. Solomon Frattini closed groups and adequate extensions of global fields

1

Leep, D. B. See H. G. Grundman

Lemańczyk, M., E. Lesigne, F. Parreau, D. Volný and M. Wierdl Random ergodic theorems and real cocycles

285

Lesigne, F. See M. Lemańczyk

Mazzeo, R. and M. Taylor Curvature and uniformization  323

Montaldo, S. See R. Caddeo Oniciuc, C. See R. Caddeo Parreau, F. See M. Lemańczyk

Rams, M. Packing dimension estimation for exceptional parameters  125

Rudolph, D. See C. Hoffman Saxl, J. See N. Gordeev Silberger, S. See E. M. Coven Smith, T. L. See H. G. Grundman Smith, T. L. See D. B. Leep Solomon, R. See D. B. Leep

Światek, G. and J. J. P. Veerman On a conjecture of Furstenberg 145

Taylor, M. See R. Mazzeo Thurston, W. See E. M. Coven Veerman, J. J. P. See G. Światek Volný, D. See M. Lemańczyk Wierdl, M. See M. Lemańczyk Zhivkov, N. V. See F. S. de Blasi